OpenStudy (anonymous):

solve the following equation x/(x+3) + 2/5 =0

OpenStudy (anonymous):

please help me out.... i really need to get this done

OpenStudy (anonymous):

x/x+3=-2/5

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

no multiply fractions by 5

OpenStudy (anonymous):

what is result?

OpenStudy (anonymous):

\[\frac{x}{x+3}+\frac{2}{5}=0\] It may help you replace X+3 with a variable, e.g. Y.

OpenStudy (anonymous):

5x=-2(x+3) 5x=-2x-6 5x+2x=-6 7x=-6

OpenStudy (anonymous):

@Pharort Please don't do the work for them. This site is meant to help people understand a subject, not to get easy answers.

OpenStudy (anonymous):

oh..! sorry, I don't know because im new in here!

OpenStudy (anonymous):

Thank you. Im not trying to get easy answers anyways.

OpenStudy (anonymous):

Glad to hear @Carlene15 :) So, do you understand how to get from the original equation to what @Pharort posted?

OpenStudy (anonymous):

i understand this subject fine... its just that this assignemnt is do in umm 4 hours and I'm in a rush to complete. So any help works:)

OpenStudy (anonymous):

well anyways, thanks